OpenStudy (anonymous):

If f(x) = x + 4, find f(3). 1 7 −1 12

OpenStudy (nurali):

f(x) = x + 4 f(3) = 3+4 f(3)=______

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

7?

OpenStudy (nurali):

Correct